# UIShift: Enhancing VLM-based GUI Agents through Self-supervised Reinforcement Learning

### v1

@@ Abstract
Training effective Vision Language Models (VLMs) for GUI agents typically relies on supervised fine-tuning (SFT) over large-scale annotated datasets, where the collection process is labor-intensive and error-prone. In this work, we propose a self-supervised inverse dynamics task to enable VLMs to learn from GUI transition pairs by inferring the action that caused that transition. This training task offers two advantages: (1) It enables VLMs to ignore variations unrelated to user actions (e.g., background refreshes, ads) and to focus on true affordances such as buttons and input fields within complex GUIs. (2) The training data can be easily obtained from existing GUI trajectories without requiring human annotation, and it can be easily scaled through automatic offline exploration. Using this training task, we propose UIShift, a framework for enhancing VLM-based GUI agents through self-supervised reinforcement learning (RL). With only 2K training samples sourced from existing datasets, two VLMs—Qwen2.5-VL-3B and Qwen2.5-VL-7B—trained with UIShift achieve competitive or superior performance on grounding tasks (ScreenSpot-series benchmarks) and GUI automation tasks (AndroidControl), compared to SFT baselines and GUI-specific models that explicitly elicit reasoning abilities during RL. Our findings suggest a potential direction for enhancing VLMs for GUI agents by leveraging more self-supervised training data in the future.
@@ 1 Introduction
Mobile GUI agents [9, 11, 20, 30, 35, 39] interpret natural language instructions and perform low-level actions (e.g., click, scroll) directly on smartphone screens. They can control diverse apps as a human would, improving accessibility for users who are visually impaired, elderly, or have their hands occupied. Breakthroughs of Vision Language Models (VLMs) [2, 4, 26, 36] have reshaped the design paradigm of mobile GUI agents, transitioning from handcrafted heuristics to learned, vision-grounded policies. However, when facing complex multi-step tasks, these models struggle to deliver satisfactory accuracy [6, 20, 21, 40]. A common approach to enhance these VLMs is to apply supervised fine-tuning (SFT) on annotated datasets, e.g., using pre-recorded GUI trajectories with human-annotated task instructions [14, 22]. Although effective, collecting task-annotated GUI trajectories remains labor-intensive and error-prone [7, 8, 14, 22]. For example, collecting the AndroidControl [14] dataset takes one year of paid annotation effort to produce 15,283 task demonstrations. The high cost of the data collection process makes this paradigm difficult to scale.
@@ 2.1 Mobile GUI Agents
Recent progress in mobile GUI agents has been driven by VLMs trained via SFT on large-scale datasets. These models learn to map instructions to GUI actions using instruction-following tasks, making high-quality annotations essential. Despite the availability of diverse GUI datasets [7, 8, 14, 15, 18, 22, 28], the quantity of high-quality annotations remains insufficient for robust training and usually requires significant human effort to scale. Prior SFT pipelines often incorporate out-of-domain image-caption pairs [11, 29] and supplement training with web and desktop data to improve cross-platform generalization [5]. As a result, the overall scale of training data tends to be large: UI-TARS [20] trains on 50B tokens; Uground [9] uses 1.3M screenshots to train a visual grounding model; OS-Atlas [32] leverages 13M UI elements for grounding pretraining. MobileVLM [31] incorporates an action prediction task during pretraining; however, but limits it to one-step transitions and still relies on annotation-heavy fine-tuning for downstream generalization. In this study, we propose a self-supervised training paradigm that eliminates reliance on large-scale human-annotated data. By constructing a -step inverse dynamics task where VLMs infer the action leading from the current to a future UI state, our approach captures long-horizon GUI dynamics and enables scalable training on large, underutilized, and unlabeled GUI transition datasets.
@@ 4.1 Experimental Setup
Models and training configurations. Using the training pipeline proposed in §3 and the open-source RL framework, VLM-R1 [25], we apply full-parameter reinforcement fine-tuning to Qwen2.5-VL-3B/7B [2] on 8A800 GPUs for 8 epochs. Our best model, UIShift-7B-2K, completes training in 26 hours with 2K UI Transition samples. Additional hyperparameter details are listed in Appendix A.
@@
Reasoning configurations and comparative setup. To study the effectiveness of chain-of-thought reasoning in VLMs, we experiment with three different configurations during training and inference: (1) both training and inference with reasoning prompts (fully reasoning-dependent); (2) training with reasoning but inference without it (inference reasoning-free); (3) training and inference both without reasoning (fully reasoning-free). To further assess the effectiveness of RL, we also train SFT baselines on the same three datasets—UI Transition, Task Automation Low, Task Automation High—under the fully reasoning-free setting.
@@ 5 Conclusion
Collecting task-annotated GUI trajectories is costly and error-prone, which constrains the scalability of supervised fine-tuning on instructional data. In this work, we propose UIShift, a self-supervised, one-stage RL framework that fine-tunes GUI-specific VLMs on the -step UI Transition task using GRPO. By predicting the first action that moves UI state to given the pair (, ), the model learns to map screen-level visual differences to low-level actions. This objective lets it generalizes to GUI grounding and task automation without extra labels, eliminates annotation costs, and unlocks large-scale, unlabeled GUI transition datasets. Without reasoning prompts at either training or inference time, UIShift-7B-2K outperforms annotation-dependent baselines on five static benchmarks, showing that UI Transition provides a strong learning signal for GUI task generalization.
This study has two main limitations. First, we train on 1K/2K samples due to compute constraints; accessing how performance scales with larger datasets is the next step. Second, our evaluation is confined to static benchmarks; future work will extend to dynamic ones such as AndroidWorld [21].

### v2

@@ Abstract
Training effective Vision Language Models (VLMs) for GUI agents typically relies on supervised fine-tuning (SFT) over large-scale annotated datasets, where the collection process is labor-intensive and error-prone. In this work, we propose a self-supervised inverse dynamics task to enable VLMs to learn from GUI transition pairs by inferring the action that caused that transition. This training task offers two advantages: (1) It enables VLMs to ignore variations unrelated to user actions (e.g., background refreshes, ads) and to focus on true affordances such as buttons and input fields within complex GUIs. (2) The training data can be easily obtained from existing GUI trajectories without requiring human annotation, and it can be easily scaled through automatic offline exploration. Using this training task, we propose UIShift, a framework for enhancing VLM-based GUI agents through self-supervised reinforcement learning (RL). With only 2K training samples sourced from existing datasets, two VLMs—Qwen2.5-VL-3B and Qwen2.5-VL-7B—trained with UIShift achieve competitive or superior performance on grounding tasks (ScreenSpot-series benchmarks) and GUI automation tasks (AndroidControl), compared to SFT baselines and GUI-specific models that explicitly elicit reasoning abilities during RL. Our findings suggest a potential direction for enhancing VLMs for GUI agents by leveraging more self-supervised training data in the future. Code, model, and data are available at: https://github.com/UbiquitousLearning/UIShift.
@@ 1 Introduction
Mobile GUI agents [9, 11, 23, 32, 37, 41] interpret natural language instructions and perform low-level actions (e.g., click, scroll) directly on smartphone screens. They can control diverse apps as a human would, improving accessibility for users who are visually impaired, elderly, or have their hands occupied. Breakthroughs of Vision Language Models (VLMs) [2, 4, 17, 22, 38] have reshaped the design paradigm of mobile GUI agents, transitioning from handcrafted heuristics to learned, vision-grounded policies. However, when facing complex multi-step tasks, these models struggle to deliver satisfactory accuracy [6, 23, 24, 42]. A common approach to enhance these VLMs is to apply supervised fine-tuning (SFT) on annotated datasets, e.g., using pre-recorded GUI trajectories with human-annotated task instructions [15, 25]. Although effective, collecting task-annotated GUI trajectories remains labor-intensive and error-prone [7, 8, 15, 25]. For example, collecting the AndroidControl [15] dataset takes one year of paid annotation effort to produce 15,283 task demonstrations. The high cost of the data collection process makes this paradigm difficult to scale.
@@ 2.1 Mobile GUI Agents
Recent progress in mobile GUI agents has been driven by VLMs trained via SFT on large-scale datasets. These models learn to map instructions to GUI actions using instruction-following tasks, making high-quality annotations essential. Despite the availability of diverse GUI datasets [7, 8, 15, 16, 20, 25, 30], the quantity of high-quality annotations remains insufficient for robust training and usually requires significant human effort to scale. Prior SFT pipelines often incorporate out-of-domain image-caption pairs [11, 31] and supplement training with web and desktop data to improve cross-platform generalization [5]. As a result, the overall scale of training data tends to be large: UI-TARS [23] trains on 50B tokens; Uground [9] uses 1.3M screenshots to train a visual grounding model; OS-Atlas [34] leverages 13M UI elements for grounding pretraining; ScaleTrack [12] uses 7.5M screenshots for grounding and planning, and introduces a backtracking task that predicts prior actions based on the current state and task instruction. MobileVLM [33] incorporates an action prediction task during pretraining, but limits it to one-step transitions and still relies on annotation-heavy fine-tuning for downstream generalization. While these approaches improve model performance, they remain dependent on large-scale annotated data and explicit instruction supervision. In this study, we propose a self-supervised training paradigm that eliminates reliance on large-scale human-annotated data. By constructing a -step inverse dynamics task where VLMs infer the action leading from the current to a future UI state, our approach captures long-horizon GUI dynamics and enables scalable training on large, underutilized, and unlabeled GUI transition datasets.
@@ 4.1 Experimental Setup
Models and training configurations. Using the training pipeline proposed in §3 and the open-source RL framework, VLM-R1 [28], we apply full-parameter reinforcement fine-tuning to Qwen2.5-VL-3B/7B [2] on 8A800 GPUs for 8 epochs. Our best model, UIShift-7B-2K, completes training in 26 hours with 2K UI Transition samples. Additional hyperparameter details are listed in Appendix A. Training prompts are available in Appendix C.1
@@
Reasoning configurations and comparative setup. To study the effectiveness of chain-of-thought reasoning in VLMs, we experiment with three different configurations during training and inference: (1) both training and inference with reasoning prompts (fully reasoning-dependent); (2) training with reasoning but inference without it (inference reasoning-free); (3) training and inference both without reasoning (fully reasoning-free). To further assess the effectiveness of RL, we also train SFT baselines on the same three datasets—UI Transition, Task Automation Low, Task Automation High—under the fully reasoning-free setting. We use a shared set of evaluation prompts for both Qwen2.5-VL baselines and our UIShift models, detailed in Appendix C.2. Scores of other methods are taken directly from the original papers.
@@ 5 Conclusion
Collecting task-annotated GUI trajectories is costly and error-prone, which constrains the scalability of supervised fine-tuning on instructional data. In this work, we propose UIShift, a self-supervised, one-stage RL framework that fine-tunes GUI-specific VLMs on the -step UI Transition task using GRPO. By predicting the first action that moves UI state to given the pair (, ), the model learns to map screen-level visual differences to low-level actions. This objective lets it generalizes to GUI grounding and task automation without extra labels, eliminates annotation costs, and unlocks large-scale, unlabeled GUI transition datasets. Without reasoning prompts at either training or inference, UIShift-7B-2K outperforms annotation-dependent baselines on five static benchmarks, showing that UI Transition provides a strong learning signal for GUI task generalization.